# Diagnosing Vision Language Models’ Perception by Leveraging Human Methods for Color Vision Deficiencies

### v1

@@ 3.3 RQ1: Generation Level
Table 4 shows model accuracy across different conditions and prompt types. Appendix D provides actual digit outputs from each model. In the Base condition, GPT-4o reaches 95 % on Normal, mPLUG-Owl3 achieves 71 %, and other models remain below 53 %. Performance collapses for both Protanopia and Deuteranopia, with no model exceeding 24 %. In contrast, Tritanopia scores exceed 65 % for GPT-4o and mPLUG-Owl3, as many plates appear visually identical to Normal. With Linguistic Support, brief descriptions shift mean accuracy by at most two points without changing rankings. GPT-4o rises to 23 % on Deuteranopia but consistently drops similarly on Tritanopia. Qwen2.5-VL often refuses when disability terms appear, averaging 0 %. Under Visual Support, example images simulating each deficiency are largely ineffective. GPT-4o drops to 41 %, Qwen2.5-VL falls by 9 %. mPLUG-Owl3 shows a modest gain on Deuteranopia, but drops on Normal and Tritanopia offset the benefit.
At the generation level, current LVLMs assume normal color vision and struggle to simulate how CVD users perceive color. Even with support, they fail to generate outputs suited to CVDs perception.
@@ 3.4 RQ2: Confidence Level
We measure token level perplexity for each forced-decoded output to evaluate how confidently LVLMs produce responses [^2] See Appendix E for details. Lower perplexity indicates higher model confidence.. Comparing perplexity across conditions reveals whether confidence aligns with accuracy, whether linguistic or visual cues reduce uncertainty, and how strongly models default to normal color vision. Figure 2 shows perplexity for mPLUG-Owl3 across four vision types and three prompts; results for other models appear in Appendix F. Perplexity remains low for Normal and Tritanopia but increases for Protanopia and Deuteranopia, suggesting that the model responds as if seeing the plates with normal color vision. Linguistic cues have little effect, and while visual support lowers perplexity for red and green deficiencies, it does not improve accuracy (Table 4).
At the confidence level, current LVLMs show consistently uncertainty in red and green tasks and fail to produce responses with appropriate confidence, even with linguistic or visual cues.
Type Prompt
Doctor-Style Inference When a person sees “{task_number}” in this Ishihara test plate, which types of color vision are possible?
Choose all that apply from the following options.
Options: - Normal - Protanopia - Deuteranopia - Tritanopia
Please output your answer.
Table 5: Doctor-style prompt, with {task_number} replaced by the correct answer for each Ishihara plate.
@@
Table 6: Color vision task accuracy (%) under Normal and Protanopia conditions for various LVLMs.
@@ 3.5 RQ3: Internal Representation Level
We analyze whether LVLMs can perform doctor-style diagnosis by inferring color vision types from Ishihara responses, using the instruction in Table 5. Doctor-style setting evaluates whether models can identify CVD types from response, as a doctor interprets test results. For each plate, we force-decode four labels, Normal, Tritanopia, Protanopia, Deuteranopia, and select the most probable. As shown in Table 6, models perform well on Protanopia and Deuteranopia (over 70 %, except Phi-3.5), but poorly on Normal and Tritanopia (under 30 %). This contrasts with Table 4, where they failed to recognize digits under red and green conditions.
Figure 3: LLaVA LogitLens averages. Left: Normal/Tritanopia; Right: Protanopia/Deuteranopia.
@@
To further examine this behavior, we apply Logit Lens nostalgebraist (2020) to analyze token level probabilities across layers, following prior work showing that intermediate signals can influence outputs Wendler et al. (2024); Schut et al. (2025). Details of this computation appear in Appendix G. Figure 3 shows average diagnosis probabilities in LLaVA-NeXT. Results for other models are in Appendix H. LLaVA-NeXT tends to interpret images as Normal in mid-layers, shifting to Protanopia at the final layer, suggesting inconsistent processing. Nearly identical curves across ground-truths imply poor differentiation between conditions.
At the internal representation level, current LVLMs do not clearly distinguish color vision types. They rely on surface cues like readable digits, rather than using perceptual reasoning.
@@ 4 Conclusion
In this study, we evaluated whether LVLMs can handle diverse aspects of human color perception, especially CVDs, by using the widely adopted Ishihara Test as a benchmark for perceptual simulation. We systematically analyzed this capability by examining generation, confidence, and internal representation. The results show that although current LVLMs have accurate linguistic knowledge about CVDs, they do not understand how CVD users perceive color. They generate inappropriate outputs, exhibit low confidence in red and green tasks, and lack internal representations that distinguish vision types, indicating a limited consideration of perceptual aspects of color. These findings highlight the need to design LVLMs that go beyond text descriptions and account for perceptual differences, making AI more accessible and user-friendly.
@@ 5 Limitations
Our evaluation relies on a small set of Ishihara plates that primarily target red and green color-vision deficiencies. The limited number and narrow focus reduce the diversity of conditions tested, so our findings may not generalise to other CVD types or broader visual tasks. Nevertheless, this short paper fulfills its objective by demonstrating the importance of analyzing model behavior in depth, specifically from the perspective of color perception, as a step toward building user-friendly and perceptually inclusive LVLMs. While scaling to broader datasets and real-world tasks remains important future work, our focus here is on establishing a methodology and highlighting key limitations in current models.
@@
The Ishihara Test is primarily designed to detect red and green deficiencies (Protanopia and Deuteranopia), and is not suitable for evaluating blue–yellow deficiencies such as Tritanopia. As such, our findings may not fully capture how LVLMs handle Tritanopia cases. Nevertheless, we included Tritanopia-specific instructions to test whether models follow the prompt condition or respond based on the visible digit. Future work should evaluate all major CVD types, including Tritanopia, using more general tasks beyond Ishihara plates.
@@
We prompt LVLMs to simulate color-blind perception in a controlled, diagnostic task (identifying digits in Ishihara patterns). This single evaluation format may not capture the full complexity of color reasoning. Prior studies show that changing task formulations alone can yield different outcomes. Moreover, we assume the Ishihara ground-truth digits as an oracle for correct behavior, without validating outputs with people who have CVD or ophthalmologists.
@@
Success or failure on stylised Ishihara plates does not guarantee that a model truly perceives color differences across diverse, real-world settings. An LVLM might exploit dot-pattern heuristics or memorised cues rather than authentically experiencing color confusion. Prior work shows vision–language models can be biased toward certain colors and struggle with others Raj et al. (2024), suggesting performance in one controlled test may not translate to broad color understanding. Caution is therefore needed when extrapolating our results to natural images with varied palettes or complex contexts.
@@ 6 Ethical Considerations
Recent AI research communities, including ACL, have emphasized the importance of designing fair and inclusive systems that accommodate users with perceptual or cognitive differences. Color-vision deficiency, which affects a significant portion of the global population, is one such factor that is often overlooked in model development and evaluation. By explicitly testing LVLM behavior under CVD conditions, this study contributes to a broader understanding of accessibility in multimodal AI systems and encourages further research in this direction. Building on this motivation, we also discuss fairness-related concerns in the context of model evaluation.
@@
The Ishihara Test Ishihara (1917) is a long-established diagnostic tool widely used in academic research across fields such as ophthalmology, psychology, education, and machine learning (Zhao et al., 2019; Klinke et al., 2024; Lin et al., 2019). We did not reproduce, modify, or redistribute the plates; they were used solely for research and analysis.
@@
Portions of the manuscript, such as prompt templates and wording adjustments, were drafted with the assistance of GPT-4 to ensure linguistic clarity. All technical content, analysis scripts, and final decisions were made by the research team.
@@ Appendix A Detailed Model Settings
In this study, to ensure fair and consistent performance comparisons, all experiments were conducted using a single NVIDIA RTX 6000 Ada GPU. All models were run under the same computational settings, with generation performed using half quantization to reduce memory usage while maintaining model fidelity. We used publicly available models from Hugging Face and OpenAI, as summarized in Table 7.
@@
Qwen2.5-VL 7B Qwen/Qwen-2.5-VL-7B-Instruc
@@
Appendix B Prompt Details
To evaluate how LVLMs respond under different levels of guidance, we design three prompt types for each color vision condition: Base, Linguistic Support, and Visual Support. Each targets a distinct aspect of the model’s capacity to simulate color perception under CVDs.
Base Prompts
Base prompts provide only the color vision condition, for example “You are Protanopic”, and ask the model to report the number shown in the Ishihara plate. This setup tests whether the model can simulate CVDs using only the condition label. See Table 8.
Prompts with Linguistic Support
These prompts add a textual description of the impairment, such as noting that red tones appear darker. This setting assesses whether linguistic context improves predictions. See Table 10.
Prompts with Visual Support
These prompts include a brief condition description and a reference example showing what number appear in another plate. The model is then asked to identify the number in a new image. This tests whether few-shot visual examples improve simulation. See Table 9. All prompts end with: “Only provide the answer. If you see nothing, answer ‘Nothing’. No additional information.” This standardization allows consistent, quantitative comparison across all settings.
Appendix C Language Knowledge of the Ishihara Test and CVDs

### v2

@@ 4.1 RQ1: Generation Level
We use Digit Accuracy to evaluate how often the model correctly identifies the numeral on each Ishihara plate. Table 4 shows accuracy across conditions and prompt types. In the Normal condition, humans are expected to score 100 %, but only GPT-4o approaches this with 90.5 %, followed by mPLUG-Owl3 at 71 %, while some models score near 0 %. Performance collapses for both Protanopia and Deuteranopia, with no model exceeding 24 %, clearly indicating the difficulty of simulating CVDs. With Linguistic Support, brief descriptions shift accuracy by at most two points without changing rankings; GPT-4o improves to 23 % on Deuteranopia but drops on Tritanopia. Qwen2.5-VL often refuses under disability-related terminology, averaging 0 %. With Visual Support, few-shot examples have minimal effect; accuracy shows no consistent improvement and sometimes decreases.
Table 5 lists the gold digits and Base-prompt predictions across 25 plates. GPT-4o reproduces most digits under Normal vision, and mPLUG-Owl3 performs well, but the remaining models are inconsistent. Under Protanopia and Deuteranopia, most models repeat Normal outputs, return empty strings, or collapse to a single digit, e.g., “70.” Thus, strong Normal accuracy does not imply successful CVD simulation. Even with Linguistic or Visual Support, accuracy improves little and sometimes introduces new errors, indicating that prompt modifications alone cannot reproduce perceptual variation. Tritanopia shares the same ground-truth digits as Normal and therefore serves as a control condition. Models output nearly identical digits for both, matching clinical answers but indicating fallback to Normal rather than simulation. Severe failures under Protanopia and Deuteranopia further show that agreement in Tritanopia does not imply perceptual simulation.
At the generation level, current LVLMs assume normal color vision and, even with prompt support, fail to generate text that matches CVDs perception.
LVLM Size Prompt Normal Protanopia Deuteranopia Tritanopia
Mean SD Mean SD Mean SD Mean SD
mPLUG‑Owl3 7B Base 3.8 6.7 69.7 78.8 69.0 116.0 3.7 6.3
Linguistic 3.8 6.7 78.9 101.5 105.9 185.6 4.2 6.9
Visual 16.0 40.2 36.5 43.3 37.1 59.4 14.3 33.5
Phi‑3.5 4.2B Base 14.5 13.3 140.5 325.5 80.0 256.1 23.9 28.6
Linguistic 14.5 13.3 334.0 792.7 158.2 544.1 24.9 28.7
Visual 7.0 6.2 13.1 12.3 12.7 12.8 7.6 7.3
Qwen2.5‑VL 7B Base 46.0 83.3 72.8 62.1 57.2 69.1 27.4 36.6
Linguistic 46.0 83.3 62.0 68.7 51.4 59.7 27.7 36.6
Visual 45.0 56.5 313.7 381.8 131.7 172.0 78.0 82.1
Table 6: Mean and standard deviation (SD) of per-token perplexity across four simulated vision conditions (Normal, Protanopia, Deuteranopia, Tritanopia) under three prompt settings (Base, Linguistic, Visual) for each LVLM. Only models that could be tested under all three settings are shown.
Figure 2: Violin plots of mPLUG-Owl3 perplexity on Ishihara digit prediction. Columns indicate the condition specified in the prompt (Normal, Protanopia, Deuteranopia, Tritanopia). Within each column, colors denote the vision type that defines the candidate digit whose perplexity is measured (Normal: blue, Protanopia: red, Deuteranopia: green, Tritanopia: yellow). Lower perplexity (for the gold answer) indicates higher model confidence.
@@ 4.2 RQ2: Confidence Level
To evaluate model confidence, we measure per-token perplexity for each forced-decoded output. Comparing perplexity across conditions reveals whether confidence aligns with accuracy, whether linguistic or visual cues reduce uncertainty, and how strongly models default to normal vision.
Table 6 reports the mean and standard deviation (SD) of perplexity (shown as mean $\pm$ SD) for all models under the three prompt settings. Across the board, Normal and Tritanopia remain lower than Protanopia and Deuteranopia, indicating a pronounced bias toward normal vision, though exceptions are observed in Llama-3.2. One model rises from about $4\pm 7$ on Normal to $70\pm 80$ on Protanopia, while another climbs from about $14\pm 13$ to $140\pm 330$. Perplexity generally increases under Linguistic and Visual Support. Although visual cues lower perplexity for a few models, the tendency persists: Normal is always lowest and the red and green deficiencies remain highest.
@@
Table 7: Doctor-style diagnosis accuracy (%) under Normal and Protanopia conditions for various LVLMs.
Figure 3: Layer-wise diagnosis probabilities for LLaVA, averaged over all plates.
Normal Protanopia Deuteranopia Tritanopia Avg.
MedVLM-R1 57.1 57.1 0.0 5.9 5.9 0.0 5.9 5.9 0.0 61.9 61.9 0.0 32.7 32.7 0.0
Qwen‑DotPattern 71.4 71.4 23.8 5.9 5.9 5.9 5.9 5.9 17.7 66.7 66.6 23.8 37.5 37.5 17.8
Table 8: Digit accuracy (%) of each LVLM under simulated vision types: Normal, Protanopia, Deuteranopia, and Tritanopia, with support settings: Base, Ling., and Vis. As for Vis., it was used only for models supporting multi-image input. This table extends Table 4 by including additional results for MedVLM-R1 and Qwen-DotPattern.
Figure 2 plots perplexity for mPLUG-Owl3. We selected this model because it supports multiple images and exhibits the most stable perplexity across prompt settings. The three prompt settings share nearly identical shapes; Normal clusters near zero, whereas Protanopia and Deuteranopia rise to tall peaks. Adding a brief description or a reference image leaves the distribution almost unchanged, which shows that prompt tweaks do not alter the model’s internal uncertainty. Other results are in Appendix E, showing the same trend: low perplexity for Normal vision, higher and variable for CVDs, with little effect from prompt changes.
At the confidence level, Normal vision consistently yields the lowest perplexity, while the CVD conditions result in similarly high perplexity, suggesting that models are fundamentally confused when simulating altered color vision and fail to reproduce such perceptual states.
@@ 4.3 RQ3: Internal Representation Level
We analyze whether LVLMs can perform doctor-style diagnosis by inferring vision types from responses using the Doctor-Style prompt (Table 3). This setting evaluates whether models can identify vision types from a response, analogous to how a doctor interprets Ishihara results. For each plate, we force-decode four labels (Normal, Protanopia, Deuteranopia, and Tritanopia) and select the most probable one. As shown in Table 7, models perform well on Protanopia (over 70%, except Phi-3.5), but poorly on Normal (under 30%). This contrasts with Table 4, where they failed to recognize digits under red and green conditions.
LVLM Size Prompt Normal Protanopia Deuteranopia Tritanopia
Mean SD Mean SD Mean SD Mean SD
Qwen2.5‑VL 7B Base 46.0 83.3 72.8 62.1 57.2 69.1 27.4 36.6
Linguistic 46.0 83.3 62.0 68.7 51.4 59.7 27.7 36.6
Visual 45.0 56.5 313.7 381.8 131.7 172.0 78.0 82.1
MedVLM-R1 2B Base 15.5 28.6 166.4 325.3 86.1 160.8 14.7 26.2
Linguistic 15.5 28.6 189.7 369.6 91.5 195.2 13.2 24.5
Visual 48.6 63.6 83.7 92.3 72.7 89.6 50.9 62.1
Qwen‑DotPattern 2B Base 4.7 8.9 79.2 136.3 75.4 171.7 4.7 6.6
Linguistic 4.7 8.9 106.3 213.0 100.2 249.9 4.7 6.5
Visual 6.2 9.1 41.5 55.2 28.7 47.1 7.4 9.9
Table 9: Mean and standard deviation (SD) of per-token perplexity across four simulated vision conditions under three prompt settings (Base, Linguistic, Visual). This table extends Table 6 by including additional results for MedVLM-R1 and Qwen-DotPattern.
LVLM Size Normal Protanopia
Qwen2.5-VL 7B 4.0 86.2
MedVLM-R1 7B 21.4 88.0
Qwen-DotPattern 7B 85.7 0.0
Table 10: Doctor-style diagnosis accuracy (%) under Normal and Protanopia conditions for MedVLM-R1 and Qwen-DotPattern, extending Table 7.
@@
To further examine this behavior, we use Layer-Wise Diagnosis Probability, inspired by Logit Lens nostalgebraist2020logitlens, to track token-level probabilities across transformer layers. Figure 3 uses LLaVA-NeXT as a representative example. Results for other models are provided in Appendix F, where no consistent trends were observed across models, indicating that the layer-wise behavior varies depending on the architecture. LLaVA-NeXT tends to interpret images as Normal in mid-layers, shifting to Protanopia at the final layer, suggesting inconsistent processing. Nearly identical curves across ground-truths imply poor differentiation between conditions.
At the internal representation level, the model exhibits nearly identical behavior across all color vision types, showing no evidence of reasoning that reflects differences between conditions. This indicates a fundamental inability to understand and reproduce variations in color perception.
4.4 Domain Knowledge and Pattern Memorization
To deepen our analysis, we conducted two focused experiments as outlined below.
Effect of medical expertise
As Ishihara plates originate in clinical practice, we expect that an LVLM trained on medical images and reports would outperform standard models. We therefore evaluated MedVLM‑R1 pan2025medvlmr1incentivizingmedicalreasoning, which is built on Qwen‑2.0‑VL wang2024qwen2 and fine‑tuned on medical data, to investigate whether training on clinical data improves performance on color-based diagnostic tools.
Dot pattern memorization hypothesis
A potential concern is that models may solve the task by exploiting superficial dot-pattern regularities in Ishihara plates, rather than demonstrating perceptual reasoning. To directly probe this, we fine‑tuned Qwen2.0‑VL on 10,000 synthetic Ishihara‑style images released by grover2025huemanityprobingfinegrainedvisual, producing a model specialized in dot‑pattern recognition. We refer to this model as Qwen‑DotPattern. This setup allows us to assess whether pattern memorization alone can explain model performance. Training details are provided in Appendix G.
Table 8 shows generation-level accuracy for MedVLM-R1 and Qwen-DotPattern. Under the Normal condition with the Base prompt, both outperform the Qwen2.5-VL baseline (MedVLM-R1: 57.1%, Qwen-DotPattern: 71.4%), indicating that domain-specific training with dot-pattern stimuli enhances recognition under typical vision. However, under Protanopia and Deuteranopia, follow the same trend as other models, showing low accuracy across prompts, suggesting that neither medical knowledge nor pattern memorization reproduces CVD perceptual distortions.
Table 9 presents the confidence level results in terms of perplexity. Under the Normal condition, both MedVLM-R1 and Qwen-DotPattern show lower perplexity than Qwen2.5-VL, indicating that domain adaptation enhances confidence on typical inputs. However, under red–green CVDs, both show higher perplexity than the base model, suggesting that medical or synthetic training improves confidence in normal vision but increases miscalibration when color perception is impaired.
We extend the doctor diagnosis analysis by adding MedVLM-R1 and Qwen-DotPattern, as shown in Table 10. MedVLM-R1 shows similar accuracy in both Normal (21.4%) and Protanopia (88.0%) settings, consistent with general model trends. This suggests that while medical training supports classification under impaired vision, it does not improve Normal vision. Qwen-DotPattern shows the opposite pattern. It performs well under Normal (85.7%) but fails under Protanopia (0.0 %). This suggests the model may rely on memorized dot layouts rather than learning generalizable structure, improving recognition in standard settings but reducing adaptability under altered vision.
These results show that training on domain-specific and dot-pattern images can improve model behavior in normal color vision, but do not lead to gains in reproducing the perceptual experience of CVDs. This highlights the value of the Ishihara Test, which remains a strong diagnostic benchmark despite its small size, as it is unaffected by pattern memorization and reliably evaluates a model’s reasoning under impaired color perception.
@@ 5 Conclusion
This work examined whether LVLMs can simulate variation in human color perception using the Ishihara Test. We analyzed model behavior at the levels of generation, confidence, and internal representation to assess whether linguistic knowledge about CVDs reflects perceptual grounding. Although LVLMs can describe CVDs and the test procedure, they default to a normative color percept and do not reproduce the digit responses associated with different vision types. Additional textual guidance or visual manipulation does not induce alternative percepts, and probing reveals only weak separation between vision types in internal activations. These findings indicate that current LVLMs lack mechanisms for alternative perceptual experiences, highlighting the need for approaches that support perceptual diversity in future multimodal systems for accessible and human-centered deployment.
@@ 6 Limitations
Our evaluation relies on a small set of Ishihara plates that primarily target red–green color-vision deficiencies. The limited number and narrow focus reduce the diversity of conditions tested, so our findings may not generalise to other CVD types or broader visual tasks. While scaling to broader datasets and real-world tasks remains important future work, our focus here is on establishing a methodology and analysing behavioural differences across prompted vision conditions rather than memorisation. Possible pre-training exposure to Ishihara plates would mainly affect the Normal condition and does not account for failures under CVD prompts.
@@
The Ishihara Test is primarily designed to detect red–green deficiencies (Protanopia and Deuteranopia) and is not suitable for assessing blue–yellow deficiencies such as Tritanopia. Moreover, the medically defined ground-truth answers for Tritanopia coincide with those for Normal vision. In our setting, Tritanopia therefore functions largely as a control for instruction following rather than as a genuine benchmark for blue–yellow deficiencies. As a result, conclusions on Tritanopia are limited, and future work should evaluate all major CVD types using more general tasks beyond Ishihara plates.
@@
We prompt LVLMs to simulate color-blind perception in a controlled diagnostic setting, tasking them with identifying digits in Ishihara plates. We use a restricted set of prompts (role-playing, doctor-perspective, step-by-step). While failures were broadly consistent across them, the prompt space is incomplete and other formulations may elicit different behaviours. Consequently, our conclusions characterise LVLM behaviour under this evaluation format rather than ruling out prompts that could induce more faithful CVD simulation. We also treat medically defined Ishihara digits as ground truth without validation from individuals with CVD or clinicians, which may limit ecological validity.
@@
Performance on stylised Ishihara plates does not imply that a model can handle color perception in broader settings. Models may rely on dot-pattern heuristics or memorised cues rather than simulating color confusion. Prior work also shows that vision–language models can exhibit color biases raj-etal-2024-biasdora, suggesting that controlled diagnostic performance may not translate to broader color understanding. Caution is therefore required when extrapolating to natural images. Finally, this study is diagnostic rather than prescriptive: we do not test mitigation strategies, and developing such methods remains an important direction for CVD-aware LVLMs.
@@ 7 Ethical Considerations
Recent AI research communities, including ACL, have emphasized the importance of designing fair and inclusive systems that accommodate users with perceptual or cognitive differences. Color vision deficiency, which affects a significant portion of the global population, is one such factor that is often overlooked in model development and evaluation. By explicitly testing LVLM behavior under CVD conditions, this study contributes to a broader understanding of accessibility in multimodal AI systems and encourages further research in this direction. Building on this motivation, we also discuss fairness-related concerns in the context of model evaluation.
@@
The Ishihara Test ishihara1917test is a long-established diagnostic tool widely used in academic research across fields such as ophthalmology, psychology, education, and machine learning (zhao2019comparison; klinke2024comparative; s19102250). We did not reproduce, modify, or redistribute the plates; they were used solely for research and analysis.
@@
Portions of the manuscript, such as prompt templates and wording adjustments, were drafted with the assistance of GPT-4 to ensure linguistic clarity. All technical content, analysis scripts, and final decisions were made by us.
@@ Appendix A Detailed Model Settings
In this study, to ensure fair and consistent performance comparisons, all experiments were conducted using a single NVIDIA RTX 6000 Ada GPU. All models were run under the same computational settings, with generation performed using half quantization to reduce memory usage while maintaining model fidelity. We used publicly available models from Hugging Face and OpenAI, as summarized in Table 11. Generation was conducted using greedy decoding, with fixed random seeds to ensure reproducibility. All other settings adhered to the default configurations of each model implementation.
@@
Qwen2.5-VL 7B Qwen/Qwen-2.5-VL-7B-Instruct
@@
Appendix B Language Knowledge of the Ishihara Test and CVDs